OpenStudy (anonymous):

Show that the vectors 6a+8b and 9a+12b are parallel

hero (hero):

6a + 8b = 2(3a + 4b) 9a + 12b = 3(3a + 4b) They're both essentially the same graph multiplied by different constants.

hero (hero):

Same thing, different approach